OpenStudy (anonymous):

Which inference can be made from the data in the two-way table? P(shirt is red|shirt is large) = P(shirt is large) P(shirt is blue|shirt is large) = P(shirt is blue) P(shirt is medium and blue) = P(shirt is medium|shirt is blue) P(shirt is medium and blue) = P(shirt is medium) P(shirt is large|shirt is red) = P(shirt is red)

OpenStudy (anonymous):

Answer the question based on the data in the table. Shirt Color Size Large Medium Total Red 42 48 90 Blue 35 40 75 Total 77 88 165

OpenStudy (misssunshinexxoxo):

A) P( shirt is red / shirt is large) = 42 /77 P( shirt is large) = 77/ 165 A is false B) P(shirt is blue|shirt is large) = 35 / 77 P( shirt is blue) = 75/165 B is false C) P( shirt is medium and blue) = 40/165 P( shirt is medium / shirt is blue) = 40 / 75 C is false D) P( shirt is medium and blue) = 40/165 P( shirt is medium) = 88/165 E) P( shirt is large / shirt is red) = 42/90 P( shirt is red) = 90/ 165 The results prove that shirt size and shirt color are not independent.

OpenStudy (anonymous):

this doesn't belong in history

OpenStudy (misssunshinexxoxo):

Very true it belongs in math but maybe this asker didn't know

OpenStudy (anonymous):

that didnt really help miss care 2 explain

OpenStudy (misssunshinexxoxo):

The answer is The results prove that shirt size and shirt color are not independent.

OpenStudy (paki):

@bsp50150 welcome to open study... please post your question in its exact location, so that it can be guided properly... thanks

OpenStudy (misssunshinexxoxo):

It's alright I answered it; the asker will know for next time. Cheers everyone